# Observational signatures of Rotating compact objects in Plasma space-time

### v1

@@ Abstract
We have investigated the characteristics of shadows cast by the Kerr black hole in the presence of plasma and compared them to those of a rotating wormhole in a uniform plasma space-time for an observer at infinity. Interestingly, for the same uniform plasma density, the apparent shadow size of the rotating wormhole is always greater than that of the Kerr black hole. To further distinguish the two compact objects we studied the deflection angle and did a comparative study in the presence of the uniform and non-uniform plasma profiles. The goal of this whole exercise is to deepen our understanding of the observational phenomena of these astrophysical objects. The analysis reveals the importance of specific plasma distribution profiles, the impact of plasma on the shadow diameter, and the behavior of deflection angles in different plasma scenarios. We have calculated constraints on the plasma parameters by considering observational data and employing analytical formulations, we have calculated constraints on the plasma parameters. Our work therefore provides valuable insights into the behavior of light rays near compact objects in plasma space-time.
March 19, 2024
@@ I Introduction
Black holes (BH) and wormholes (WH) are geometrically two different astrophysical objects, however observationally the direct evidence of their existence is still not full-proof to say the least. Although both BH and WH are solutions to Einstein’s equations, they are distinctly different. Interestingly, recent shadow images of M87* Akiyama et al. (2019) and SgrA* Akiyama et al. (2022) give us hope for the existence of such objects. The wormhole remains mysterious so far, with the only similarity that exists between them being that they are both solutions of Einstein’s Field Equations. Wormholes are among the most exciting objects predicted by Einstein’s general theory of relativity, which topologically have a tunnel-like structure connecting two different regions of space-time. The first description of such a geometry appeared as early as 1916 Flamm (1916) in a study of the spatial part of the Schwarzschild metric. The idea was later refined by Einstein and Rosen, which is now known as the Einstein-Rosen bridge Einstein and Rosen (1935). However, it was more of a mathematical construction, since it did not deal with traversable WHs. Ellis Ellis (1973) first came up with the traversable general relativistic solution of the WH, and later on, Morris and Thorne Morris and Thorne (1988) studied the properties of such WHs in detail. Although WHs are hypothetical and topologically non-trivial structures in space-time, it is true that, from a theoretical standpoint, the possibility of their existence is difficult to answer. However, they should not be completely ruled out as they can very well play the role of a black hole mimickers or could be thought of as one of the many different exotic compact objects in the universe other than BHs. A compact review on the efforts to search for astrophysical wormholes (with an exhaustive list of references) can be found in Bambi and Stojkovic (2021). Theoretically, the WH can act as a lensing object similarly to the BH and the derivation of the lensing equations would lead to a similar procedure as in the case of the BH Perlick (2004). Soon after, a plethora of works have been done to distinguish the wormhole from other compact objects with the help of gravitational lensing Chetouani and Clément (1984); Clement (1984); Perlick (2004); Nandi et al. (2006); Dey and Sen (2008); Tsukamoto and Harada (2013); Kardashev et al. (2007); Cramer et al. (1995); Harko et al. (2009); Tsukamoto et al. (2012); Bambi (2013). However, a shred of concrete evidence is still lacking about the existence of wormholes in the universe. It is commonly believed that the supermassive objects at the galactic centers are black holes; however, their existence depends on the direct detection of the event horizon. Several tests have been performed to detect the event horizon Broderick et al. (2009, 2015); Narayan and McClintock (2008) but the existence of such event horizons is still questionable Abramowicz et al. (2002). Apart from this, the existence of the shadow boundary does not necessarily give evidence of the existence of the black hole since other compact objects, such as hard spheres Broderick and Narayan (2006), naked singularities Bambi and Freese (2009); Ortiz et al. (2015); Shaikh et al. (2019), non-rotating wormholes Bambi (2013); Ohgami and Sakai (2015, 2016); Mustafa et al. (2021), and rotating wormholes Shaikh (2018a), can also produce similar shadow boundaries. Such type of work including the possibility of the M87* being a superspinar within the Kerr geometry has been ruled out using the circularity of the image Bambi et al. (2019a) exists in the literature. Furthermore, after the publication of results on $*{}^{*}$ and Sagittarius $*{}^{*}$ (Sgr $*{}^{*}$) black hole shadows by Event Horizon Telescope (EHT), interest started to grow for looking into the possibility of wormhole shadows as well Ohgami and Sakai (2015); Gyulchev et al. (2018); Rahaman et al. (2021a); Ohgami and Sakai (2016). In this context, it is worth mentioning that black hole shadows in different modified gravity space-times as well as for compact objects other than black holes, such as naked singularities, hard surfaces, etc., have been studied in great depth in the past (see Perlick and Tsupko (2022); Zakharov (2014); Takahashi (2005); Cunha et al. (2015); Li and Bambi (2014); Abdujabbarov et al. (2016a); Amir and Ghosh (2016); Cunha et al. (2017); Atamurotov et al. (2013a, b); Wang et al. (2017); Papnoi et al. (2014); Uniyal et al. (2022); Abdujabbarov et al. (2015); Okyay and Övgün (2022); Uniyal et al. (2023a, b); Devi et al. (2023); Roy and Chakrabarti (2020); Wang et al. (2023); Uniyal et al. (2024); Vagnozzi et al. (2020); Roy et al. (2022); Ditta et al. (2021); Javed et al. (2022); Mustafa et al. (2022); Jusufi et al. (2019) for a few references). The work has also been done by including plasma space-time around the black hole Atamurotov and Ahmedov (2015); Abdujabbarov et al. (2016b) as well as for the Kerr de-Sitter background Chowdhuri and Bhattacharyya (2021). A lot of work is also available in the literature to understand the bending of light around the compact object (see Cramer et al. (1995); Nandi et al. (2006); Nakajima and Asada (2012); Tsukamoto et al. (2012); Tsukamoto and Harada (2017); Jusufi and Övgün (2018); Ghosh and Bhattacharyya (2022); Chowdhuri et al. (2023) for a few references) and in the presence of the plasma medium Rogers (2015a); Bisnovatyi-Kogan and Tsupko (2010a).
@@
It is well known that all astrophysical phenomena occur in a plasma-filled spacetime which eventually affects the light trajectories due to it being a dispersive medium Perlick (2000) in addition to the gravitational influence of compact objects. These plasma distributions don’t affect the observation in all frequency ranges, but they have a significant impact on low-frequency, i.e., radio frequency ranges Muhleman et al. (1970); Perlick and Tsupko (2023) in which EHT works The Event Horizon Telescope Collaboration (2019). As we know, the plasma medium has an impact on pulsar signal travel time Lorimer and Kramer (2004). To explain such behavior, one can assume that the medium is a non-magnetized pressureless plasma and that a linearized theory is sufficient to describe the gravitational field Perlick et al. (2015). It is, therefore, interesting to consider the plasma medium and look at the effect of the plasma on such radio frequency ranges. It is safe to assume that black holes as well as exotic compact objects are surrounded by a plasma in a realistic astrophysical setting, and it will therefore be an interesting approach to investigate the observable effects of the plasma, if any, on radio signals that come close to such exotic objects. In case of black holes and compact objects, studies on understanding the imaging and effect of the plasma on deflection of electromagnetic waves (with different space-time metrics) have been performed in detail by Perlick and others; see Perlick and Tsupko (2017, 2023); Tsupko and Bisnovatyi-Kogan (2013); Liu et al. (2017); Rogers (2017a, b); Er and Mao (2014); Bisnovatyi-Kogan and Tsupko (2015) and references therein.
@@
In this manuscript, our primary objective is to obtain analytical expressions for the shadow boundary of the Kerr black hole within plasma space-time, particularly for an observer positioned at infinity. We stress that although such a result may seem very trivial and possibly present in the literature, to our surprise, we found that it is not the case. A similar analytical study for high-spin Kerr has been conducted by Yan et al. Yan (2019). However, he has assumed incorrect plasma profiles that can’t satisfy the separability condition mentioned by Perlick et al. Perlick and Tsupko (2017). One of the new results discussed in this paper is the expression for the shadow boundary of the Kerr black hole surrounded by plasma with an observer situated at spatial infinity. Because most astrophysical objects are rotating, we compare the rotating BH geometry with a rotating WH geometry. The Kerr geometry is widely accepted with its uniqueness as a rotating axisymmetric solution of Einstein’s Field Equations; therefore, we compare the most simple and widely accepted solution for rotating BH, i.e. the Kerr BH with the rotating WH Teo (1998a). Some words are now in order about the rotating WH solutions proposed by Teo (1998a). The rotating Teo WH solution is based on the traversable Morris-Thorne (MT) WH Morris and Thorne (1988) solution. One of the key features in the MT WH solution is that the conditions that a traversable WH must satisfy is first listed out, and then Einstein equations are used to deduce the form of the matter required to maintain the WH, which is just the opposite of the standard procedure of using the matter content to solve the Einstein’s equations. It was also shown that the MT WH metric, together with the wormhole-shaping and traversality conditions, imply that the corresponding stress-energy tensor violates the null, and therefore also the weak, energy condition Visser (1996). Such a matter content was termed as ‘exotic’ which in turn showed that it is almost impossible to do interstellar travels using WHs. However, not everything was lost, and it was shown Visser (1989) that by giving up on spherical symmetry, it is possible for an observer falling through the wormhole to avoid encountering the exotic matter. Motivated by this, Teo Teo (1998a) has proposed the stationary, axisymmetric generalization of the MT WH. Although the null energy condition is generically violated at the throat in such a WH, but it is possible for geodesics falling through such WHs to avoid the energy-condition violating matter and thereby avoiding any problems related to that. In this work, therefore, we seek to conduct a comparative investigation between the shadows of the Kerr black hole and that of the wormhole in a homogeneous plasma space-time, provided the plasma distribution satisfies the Hamiltonian separability criterion Perlick and Tsupko (2017); Bezdekova et al. (2022); Kumar et al. (2023).
@@
The paper is organized as follows: Section II presents a brief overview of the Hamiltonian formalism for null geodesics in plasma space-time. Moving on to Section III, we derive the precise expressions for the null geodesic equations, focusing on specific plasma profiles that satisfy the crucial condition for the existence of Carter’s constant. In Section IV, our focus shifts to the celestial coordinates of the shadow boundary, where we conduct a thorough comparison of shadows for varying plasma densities of multiple plasma distributions. Additionally, this section explores the intriguing contrast between the shadows of the Kerr black hole and the Teo wormhole in a homogeneous plasma space-time. Section V delves into the calculations of the deflection angle by a slow-rotating black hole, employing the weak field approximation for both uniform and non-uniform plasma space-times. The results are compared with those obtained for the rotating wormhole Kumar et al. (2023). Lastly, Section VI is dedicated to the effort of constraining the plasma parameters, taking into account the observational data gathered from $*{}^{*}$. Throughout the paper, we have adopted units, setting $\hbar=G=c=m=1$, and employed a signature of (-,+,+,+).
@@ IV SHADOW OF KERR BLACKHOLE IN PLASMA space-time
$\xi=\left.\bigg{[}\frac{1}{\Delta}(r^{2}+a^{2}-a\eta)^{2}-\frac{\Omega_{r}}{\omega_{o}^{2}}\Bigg{]}\right|_{r=r_{ph}},$ (26)
@@ V.1 Uniform plasma distribution
$\begin{split}\Phi_{0}&=\frac{2R}{r\sqrt{r^{2}-R^{2}}},\\ \Phi_{1}&=\left(\frac{2a\left(r^{2}(R+2)-rR^{2}-R^{3}\right)+4R^{2}(r+R)}{r^{2}(r+R)\sqrt{r^{2}-R^{2}}}\right)\frac{m}{R},\\ \Phi_{2}&=\left(\frac{2(a+R)}{(r+R)\sqrt{r^{2}-R^{2}}}\right)\frac{mk_{0}}{R^{3}}.\end{split}$ (48)
@@ V.2 Non-uniform plasma distribution
$\begin{split}\psi_{0}&=\frac{2R}{r\sqrt{r^{2}-R^{2}}},\\ \psi_{1}&=\Bigg{(}\frac{4R^{2}}{r^{2}\sqrt{r^{2}-R^{2}}},\\ &\quad+a\frac{2\left(r^{2}(R+2)-rR^{2}-R^{3}\right)}{r^{2}(r+R)\sqrt{r^{2}-R^{2}}}\Bigg{)}\frac{m}{R},\\ \psi_{2}&=\Bigg{(}-\frac{1}{rR^{3}\sqrt{r^{2}-R^{2}}}\\ &\quad\left(\frac{-(a-2)r^{2}+arR-aR(R+4)}{r^{2}R^{3}(r+R)\sqrt{r^{2}-R^{2}}}\right)\frac{m}{R}\Bigg{)}\frac{k_{0}}{R^{2}}.\end{split}$ (51)
@@
$\alpha=\pm\left(\frac{4m}{R}+\frac{4ma}{R^{2}}+\left(\frac{2m}{R^{3}}-\frac{\pi}{2R^{2}}\right)k_{r}\right),$ (52)

### v2

@@ Abstract
We have investigated the characteristics of shadows cast by the Kerr black hole in the presence of plasma and compared them to those of a rotating wormhole in a uniform plasma space-time for an observer at infinity. Interestingly, for the same uniform plasma density, the apparent shadow size of the rotating wormhole is always greater than that of the Kerr black hole. To further distinguish the two compact objects we studied the deflection angle and did a comparative study in the presence of the uniform and non-uniform plasma profiles. The goal of this whole exercise is to deepen our understanding of the observational phenomena of these astrophysical objects. The analysis reveals the importance of specific plasma distribution profiles, the impact of plasma on the shadow diameter, and the behavior of deflection angles in different plasma scenarios. We have calculated constraints on the plasma parameters by considering observational data and employing analytical formulations. Our work therefore provides valuable insights into the behavior of light rays near compact objects in plasma space-time.
May 2, 2024
@@ I Introduction
Black holes (BH) and wormholes (WH) are geometrically two different astrophysical objects, however observationally the direct evidence of their existence is still not full-proof to say the least. Although both BH and WH are solutions to Einstein’s equations, they are distinctly different. Interestingly, recent shadow images of M87* Akiyama et al. (2019) and SgrA* Akiyama et al. (2022) give us hope for the existence of such objects. The wormhole remains mysterious so far, with the only similarity that exists between them being that they are both solutions of Einstein’s Field Equations. Wormholes are among the most exciting objects predicted by Einstein’s general theory of relativity, which topologically have a tunnel-like structure connecting two different regions of space-time. The first description of such a geometry appeared as early as 1916 Flamm (1916) in a study of the spatial part of the Schwarzschild metric. The idea was later refined by Einstein and Rosen, which is now known as the Einstein-Rosen bridge Einstein and Rosen (1935). However, it was more of a mathematical construction, since it did not deal with traversable WHs. Ellis Ellis (1973) first came up with the traversable general relativistic solution of the WH, and later on, Morris and Thorne Morris and Thorne (1988) studied the properties of such WHs in detail. Although WHs are hypothetical and topologically non-trivial structures in space-time, it is true that, from a theoretical standpoint, the possibility of their existence is difficult to answer. However, they should not be completely ruled out as they can very well play the role of a black hole mimickers or could be thought of as one of the many different exotic compact objects in the universe other than BHs. A compact review on the efforts to search for astrophysical wormholes (with an exhaustive list of references) can be found in Bambi and Stojkovic (2021). Theoretically, the WH can act as a lensing object similarly to the BH and the derivation of the lensing equations would lead to a similar procedure as in the case of the BH Perlick (2004). Soon after, a plethora of works have been done to distinguish the wormhole from other compact objects with the help of gravitational lensing Chetouani and Clément (1984); Clement (1984); Perlick (2004); Nandi et al. (2006); Dey and Sen (2008); Tsukamoto and Harada (2013); Kardashev et al. (2007); Cramer et al. (1995); Harko et al. (2009); Tsukamoto et al. (2012); Bambi (2013). However, a shred of concrete evidence is still lacking about the existence of wormholes in the universe. It is commonly believed that the supermassive objects at the galactic centers are black holes; however, their existence depends on the direct detection of the event horizon. Several tests have been performed to detect the event horizon Broderick et al. (2009, 2015); Narayan and McClintock (2008) but the existence of such event horizons is still questionable Abramowicz et al. (2002). Apart from this, the existence of the shadow boundary does not necessarily give evidence of the existence of the black hole since other compact objects, such as hard spheres Broderick and Narayan (2006), naked singularities Bambi and Freese (2009); Ortiz et al. (2015); Shaikh et al. (2019), non-rotating wormholes Bambi (2013); Ohgami and Sakai (2015, 2016); Mustafa et al. (2021), and rotating wormholes Shaikh (2018a), can also produce similar shadow boundaries. Such type of work including the possibility of the M87* being a superspinar within the Kerr geometry has been ruled out using the circularity of the image Bambi et al. (2019a) exists in the literature. Furthermore, after the publication of results on M87 and Sagittarius A (Sgr A) black hole shadows by Event Horizon Telescope (EHT), interest started to grow for looking into the possibility of wormhole shadows as well Ohgami and Sakai (2015); Gyulchev et al. (2018); Rahaman et al. (2021a); Ohgami and Sakai (2016). In this context, it is worth mentioning that black hole shadows in different modified gravity space-times as well as for compact objects other than black holes, such as naked singularities, hard surfaces, etc., have been studied in great depth in the past (see Perlick and Tsupko (2022); Zakharov (2014); Takahashi (2005); Cunha et al. (2015); Li and Bambi (2014); Abdujabbarov et al. (2016a); Amir and Ghosh (2016); Rosa and Rubiera-Garcia (2022); Cunha et al. (2017); Atamurotov et al. (2013a, b); Wang et al. (2017); Papnoi et al. (2014); Uniyal et al. (2022); Abdujabbarov et al. (2015); Okyay and Övgün (2022); Uniyal et al. (2023a, b); Rosa et al. (2024); Devi et al. (2023); Roy and Chakrabarti (2020); Wang et al. (2023); Uniyal et al. (2024); Vagnozzi et al. (2020); Roy et al. (2022); Ditta et al. (2021); Javed et al. (2022); Mustafa et al. (2022); Jusufi et al. (2019); Rosa et al. (2022) for a few references). The work has also been done by including plasma space-time around the black hole Atamurotov and Ahmedov (2015); Abdujabbarov et al. (2016b) as well as for the Kerr de-Sitter background Chowdhuri and Bhattacharyya (2021). A lot of work is also available in the literature to understand the bending of light around the compact object (see Cramer et al. (1995); Nandi et al. (2006); Nakajima and Asada (2012); Tsukamoto et al. (2012); Tsukamoto and Harada (2017); Jusufi and Övgün (2018); Ghosh and Bhattacharyya (2022); Chowdhuri et al. (2023) for a few references) and in the presence of the plasma medium Rogers (2015a); Bisnovatyi-Kogan and Tsupko (2010a).
@@
It is well known that all astrophysical phenomena occur in a plasma-filled spacetime which eventually affects the light trajectories due to it being a dispersive medium Perlick (2000) in addition to the gravitational influence of compact objects. These plasma distributions do not affect the observation in all frequency ranges, but they have a significant impact on low-frequency, i.e., radio frequency ranges Muhleman et al. (1970); Perlick and Tsupko (2023) in which EHT works The Event Horizon Telescope Collaboration (2019). As we know, the plasma medium has an impact on pulsar signal travel time Lorimer and Kramer (2004). To explain such behavior, one can assume that the medium is a non-magnetized pressureless plasma and that a linearized theory is sufficient to describe the gravitational field Perlick et al. (2015). It is, therefore, interesting to consider the plasma medium and look at the effect of the plasma on such radio frequency ranges. It is safe to assume that black holes as well as exotic compact objects are surrounded by a plasma in a realistic astrophysical setting, and it will therefore be an interesting approach to investigate the observable effects of the plasma, if any, on radio signals that come close to such exotic objects. In case of black holes and compact objects, studies on understanding the imaging and effect of the plasma on deflection of electromagnetic waves (with different space-time metrics) have been performed in detail by Perlick and others; see Perlick and Tsupko (2017, 2023); Tsupko and Bisnovatyi-Kogan (2013); Liu et al. (2017); Rogers (2017a, b); Er and Mao (2014); Bisnovatyi-Kogan and Tsupko (2015); Zhang et al. (2023) and references therein.
@@
In this manuscript, our primary objective is to obtain analytical expressions for the shadow boundary of the Kerr black hole within plasma space-time, particularly for an observer positioned at infinity. We stress that although such a result may seem very trivial and possibly present in the literature, to our surprise, we found that it is not the case. A similar analytical study for high-spin Kerr has been conducted by Yan et al. Yan (2019). However, he has assumed incorrect plasma profiles that can’t satisfy the separability condition mentioned by Perlick et al. Perlick and Tsupko (2017). One of the new results discussed in this paper is the expression for the shadow boundary of the Kerr black hole surrounded by plasma with an observer situated at spatial infinity. Because most astrophysical objects are rotating, we compare the rotating BH geometry with a rotating WH geometry. The Kerr geometry is widely accepted with its uniqueness as a rotating axisymmetric solution of Einstein’s Field Equations; therefore, we compare the most simple and widely accepted solution for rotating BH, i.e. the Kerr BH with the rotating WH Teo (1998a). Some words are now in order about the rotating WH solutions proposed by Teo (1998a). The rotating Teo WH solution is based on the traversable Morris-Thorne (MT) WH Morris and Thorne (1988) solution. One of the key features in the MT WH solution is that the conditions that a traversable WH must satisfy is first listed out, and then Einstein equations are used to deduce the form of the matter required to maintain the WH, which is just the opposite of the standard procedure of using the matter content to solve the Einstein’s equations. It was also shown that the MT WH metric, together with the wormhole-shaping and traversability conditions, imply that the corresponding stress-energy tensor violates the null, and therefore also the weak, energy condition Visser (1996). Such a matter content was termed as ‘exotic’ which in turn showed that it is almost impossible to do interstellar travels using WHs. However, not everything was lost, and it was shown Visser (1989) that by giving up on spherical symmetry, it is possible for an observer falling through the wormhole to avoid encountering the exotic matter. Motivated by this, Teo Teo (1998a) has proposed the stationary, axisymmetric generalization of the MT WH. Although the null energy condition is generically violated at the throat in such a WH, but it is possible for geodesics falling through such WHs to avoid the energy-condition violating matter and thereby avoiding any problems related to that. In this work, therefore, we seek to conduct a comparative investigation between the shadows of the Kerr black hole and that of the wormhole in a homogeneous plasma space-time, provided the plasma distribution satisfies the Hamiltonian separability criterion Perlick and Tsupko (2017); Bezdekova et al. (2022); Kumar et al. (2023).
@@
The paper is organized as follows: Section II presents a brief overview of the Hamiltonian formalism for null geodesics in plasma space-time. Moving on to Section III, we derive the precise expressions for the null geodesic equations, focusing on specific plasma profiles that satisfy the crucial condition for the existence of Carter’s constant. In Section IV, our focus shifts to the celestial coordinates of the shadow boundary, where we conduct a thorough comparison of shadows for varying plasma densities of multiple plasma distributions. Additionally, this section explores the intriguing contrast between the shadows of the Kerr black hole and the Teo wormhole in a homogeneous plasma space-time. Section V delves into the calculations of the deflection angle by a slow-rotating black hole, employing the weak field approximation for both uniform and non-uniform plasma space-times. The results are compared with those obtained for the rotating wormhole Kumar et al. (2023). Lastly, Section VI is dedicated to the effort of constraining the plasma parameters, taking into account the observational data gathered from M87. Throughout the paper, we have adopted units, setting $\hbar=G=c=m=1$, and employed a signature of (-,+,+,+).
@@ IV SHADOW OF KERR BLACKHOLE IN PLASMA space-time
$\xi=\left.\bigg{[}\frac{1}{\Delta}(r^{2}+a^{2}-a\eta)^{2}-\frac{\Omega_{r}}{\omega_{o}^{2}}\Bigg{]}\right|_{r=r_{c}},$ (26)
@@ V.1 Uniform plasma distribution
$\begin{split}\Phi_{0}&=\frac{R}{r\sqrt{r^{2}-R^{2}}},\\ \Phi_{1}&=\left(\frac{r^{2}(2a+R)+rR^{2}+R^{3}}{r^{2}(r+R)\sqrt{r^{2}-R^{2}}}\right)\frac{m}{R},\\ \Phi_{2}&=\left(\frac{(a+R)}{(r+R)\sqrt{r^{2}-R^{2}}}\right)\frac{mk_{0}}{R}.\end{split}$ (48)
@@ V.2 Non-uniform plasma distribution
$\begin{split}\psi_{0}&=\frac{R}{r\sqrt{r^{2}-R^{2}}},\\ \psi_{1}&=\Bigg{(}\frac{r^{2}+rR+R^{2}}{r^{2}(r+R)\sqrt{r^{2}-R^{2}}}+\\ &\quad+a\frac{2}{R(r+R)\sqrt{r^{2}-R^{2}}}\Bigg{)}m,\\ \psi_{2}&=\Bigg{(}-\frac{1}{2rR\sqrt{r^{2}-R^{2}}}-\frac{2ma}{r^{2}R(r+R)\sqrt{r^{2}-R^{2}}}+\\ &\quad\frac{r^{2}+rR-R^{2}}{2r^{2}R^{3}(r+R)\sqrt{r^{2}-R^{2}}}m\Bigg{)}k_{r}.\end{split}$ (51)
@@
$\alpha=\pm\left(\frac{4m}{R}+\frac{4ma}{R^{2}}-\left(\frac{2m}{R^{3}}+\frac{\pi}{2R^{2}}\right)k_{r}\right),$ (52)